Click here to read the eBook: Market Value Ratios P/E AND STOCK PRICE Ferrell Inc. recently reported net income of $2 million. It has 260,000 shares of common stock, which currently trades at $23 a share. Ferrell continues to expand and anticipates that 1 year from now, its net income will be $3.1 million. Over the next year, it also anticipates issuing an additional 91,000 shares of stock so that 1 year from now it will have 351,000 shares of common stock. Assuming Ferrells price/earnings ratio remains at its current level, what will be its stock price 1 year from now? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

P/E Ratio = Price per shares/Earnings per share

= $23/($2,000,000/260,000)

= 2.99

Stock price one year from now:

= P/E Ratio×Earnings per share

= 2.99×($3,100,000/351,000)

= 2.99×$8.83

= $26.41

Hence, stock in one year is $26.41